OpenStudy (anonymous):

find the equation of the perpendicular bisector of the line joining the points (-7, 5) and (-2, -6)

OpenStudy (anonymous):

slope is -11/5 the equation of a line which is perpendicular will have a slope which is the negaive reciprocal... which is 5/11. so now you have y = (5/11)x + b and you need to find b. This line is the bisector, so it crosses through the middle of the line segment, which is ((-7+-2)/2, (5+-6)/2) = (-4.5, -.5) so you plug x = -4 1/2 and y = -1/2 and solve for b